# Study of semileptonic $B\to DP\ell^{+}\nu_{\ell}$ decays based on the SU(3) flavor symmetry

### v1

@@ Abstract
Decays $B\to DP\ell^{+}\nu_{\ell}\leavevmode\nobreak\ (\ell=e,\mu,\tau)$ with the non-resonance, the charmed vector resonances, the charmed scalar resonances and the charmed tensor resonances are calculated by using the SU(3) flavor symmetry. Firstly, the decay amplitudes of different modes are related by the SU(3) flavor symmetry. Then, relevant experiential data are used to constrain nonperturbative coefficients in the non-resonant and various resonant $B\to DP\ell^{+}\nu_{\ell}$ decays. Finally, using the constrained nonperturbative coefficients, the branching ratios of not-yet-measured $B\to D^{*}P\ell^{+}\nu_{\ell}$ decays with the non-resonant and various charmed resonant contributions are predicted. Many branching ratios are predicted for the first time. We find that $B\to D\eta^{\prime}\ell^{+}\nu_{\ell},B_{s}\to D_{s}\eta^{\prime}\ell^{+}\nu_{\ell}$ decays only receive the non-resonant contributions, $B\to D_{s}K\ell^{+}\nu_{\ell}$, $B_{s}\to DK\ell^{+}\nu_{\ell}$, $B\to D\eta\ell^{+}\nu_{\ell}$ and $B_{s}\to D_{s}\eta\ell^{+}\nu_{\ell}$ decays receive both non-resonant and charmed tensor resonant contributions, $B^{+}\to D^{-}\pi^{+}\ell^{+}\nu_{\ell}$ decays receive the non-resonant, the charmed scalar resonant and the charmed tensor resonant contributions, and other $B\to D\pi\ell^{+}\nu_{\ell}$ decays receive all four kinds of contributions. These results can be used to test the SU(3) flavor symmetry approach in the four-body semileptonic decays by the future LHCb and Belle-II experiments.
@@ I Introduction
where $\mathcal{B}_{T,M}$ denote the total and $M$ resonant branching ratios. For $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ and $\mathcal{B}(B^{0}\to D^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ given in Eq. (1) and Eq. (4), the PDG reports decays via the $D_{0}$ and $D^{*}_{2}$ resonances. With the experiment running, more relevant decays could be measured at LHCb and Belle-II. Present measurements of the $B\to DP\ell^{\prime+}\nu_{\ell^{\prime}}$ decays give us an opportunity to test theoretical approaches of the $B\to DP\ell^{+}\nu_{\ell}$ decays and to predict many non-measured decays, which can be further tested by LHCb and Belle-II in the near future.
@@
Theoretically, semileptonic decays are relatively simple, since the weak and strong dynamics can be separated in these decays. All the strong dynamics in the initial and final hadrons is included in the hadronic transition form factors. The analytic structure of the $B\to DP$ form factors is more complicated than ones of the $B\to D$ or $B\to P$ form factors Feldmann:2018kqr . Due to our poor understanding of hadronic interactions, the $B\to DP$ evaluations of the form factors are difficult. In the absence of reliable calculations, the symmetry analysis can provide very useful information about the decays. SU(3) flavor symmetry is one of popular symmetry approaches. SU(3) flavor symmetry has been widely used to study hadron decays, including $b$-hadron decays He:1998rq ; He:2000ys ; Fu:2003fy ; Hsiao:2015iiu ; He:2015fwa ; He:2015fsa ; Deshpande:1994ii ; Gronau:1994rj ; Gronau:1995hm ; Shivashankara:2015cta ; Zhou:2016jkv ; Cheng:2014rfa ; Wang:2021uzi ; Wang:2020wxn , $c$-hadron decays Wang:2021uzi ; Wang:2020wxn ; Grossman:2012ry ; Pirtskhalava:2011va ; Savage:1989qr ; Savage:1991wu ; Altarelli:1975ye ; Lu:2016ogy ; Geng:2017esc ; Geng:2018plk ; Geng:2017mxn ; Geng:2019bfz ; Wang:2017azm ; Wang:2019dls ; Wang:2017gxe ; Qiao:2024srw ; Muller:2015lua , and light hadron decays Wang:2019alu ; Wang:2021uzi ; Xu:2020jfr ; Chang:2014iba ; Zenczykowski:2005cs ; Zenczykowski:2006se ; Cabibbo:1963yz . SU(3) flavor breaking effects due to the mass difference of $u,d,s$ quarks have also been studied, for instance, in Refs. Martinelli:2022xir ; Wang:2022fbk ; Imbeault:2011jz ; Wu:2005hi ; Dery:2020lbc ; Sasaki:2008ha ; Pham:2012db ; Geng:2018bow ; Flores-Mendieta:1998tfv ; Cheng:2012xb ; Xu:2013dta ; He:2014xha ; Yang:2015era .
@@
Some four-body semileptonic decays $B/D\to PP\ell^{+}\nu_{\ell}$ and $B\to D^{(*)}P\ell^{+}\nu_{\ell}$ have been studied, for instance, in Refs. Cheng:1993ah ; Tsai:2021ota ; Feldmann:2018kqr ; Kim:1999gm ; Kim:1998nn ; LeYaouanc:2018zyo ; Boer:2016iez ; Hambrock:2015aor ; Faller:2013dwa ; Shi:2021bvy ; Achasov:2020qfx ; Wiss:2007mr ; Wang:2016wpc ; Achasov:2021dvt ; Gustafson:2023lrz . In this work, we will study the $B\to DP\ell^{+}\nu_{\ell}$ decays with the non-resonant, the charmed vector resonant, the charmed scalar resonant and the charmed tensor resonant contributions by the SU(3) flavor symmetry based on the experimental data. Firstly, we will construct the hadronic amplitude relations or the form factor relations between different decay modes. Then, we will use the available data to extract the hadronic amplitudes or the form factors. Finally, we will predict the not-yet-measured modes for further tests in experiments.
@@
This paper is organized as follows. The non-resonant contributions of the $B\to DP\ell\nu_{\ell}$ decays are discussed in Sec. II. The charmed vector resonant, the charmed scalar resonant and the charmed tensor resonant contributions of $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays are presented in Sec. III. Finally, summary is given in Sec. IV.
@@ II Non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays
II.1 Meson multiplets
Since the SU(3) flavor analysis is based on the SU(3) flavor group, we give relevant meson multiplets first. Bottom pseudoscalar triplet $B_{i}$, charm pseudoscalar triplet $D_{i}$, charm scalar triplet $D_{0i}$ and charm vector triplet $D^{*}_{i}$ under the SU(3) flavor symmetry of $u$, $d$, $s$ quarks are
@@ II.1 Meson multiplets
where $i=1,2,3$ for $u$, $d$, $s$ quarks. Note that the structures of scalar $D_{0}$ mesons are not known well, and they might be four-quark states, $D_{0i}^{4q}=\Big{(}\overline{D}^{0}_{0}(\bar{c}u\bar{d}d),D^{-}_{0}(\bar{c}d\bar{u}u),D^{-}_{s0}(\bar{c}s(\bar{u}u+\bar{d}d)/\sqrt{2})\Big{)}$ Cheng:2003kg .
The charm tensor triplet $D^{*}_{2i}$ is
$D^{*}_{2i}$ $=$ $\Big{(}\overline{D}^{*}_{2}(2460)^{0},\leavevmode\nobreak\ D^{*}_{2}(2460)^{-},\leavevmode\nobreak\ D^{*}_{s2}(2573)^{-}\Big{)}.$ (10)
Note that $D_{0i}$, $D^{*}_{i}$ and $D^{*}_{2i}$ will be used for the resonances in Sec. III, and they are given here together. In addition, charm axial vector mesons $D^{\prime}_{1}/D_{1}$ have been studied in Ref. Wan:2024ssl for $B\to D^{*}P\ell^{+}\nu_{\ell}$ decays, $D^{\prime}_{1}/D_{1}$ do not contribute to $B\to DP\ell^{+}\nu_{\ell}$ decays, so relevant multiplets are not listed here. Light pseudoscalar octets and singlets $P^{i}_{j}$ is
@@
II.2 Decay amplitudes
The four-body semileptonic $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays are generated by $\bar{b}\rightarrow\bar{c}\ell^{+}\nu_{\ell}$ transition, and the effective Hamiltonian is
$\mathcal{H}_{eff}(\bar{b}\rightarrow\bar{c}\ell^{+}\nu_{\ell})=\frac{G_{F}}{\sqrt{2}}V_{cb}\bar{b}\gamma^{\mu}(1-\gamma_{5})\bar{c}\leavevmode\nobreak\ \bar{\nu_{\ell}}\gamma_{\mu}(1-\gamma_{5})\ell,$ (22)
where $G_{F}$ is the Fermi constant, and $V_{cb}$ is the CKM matrix element. Decay amplitudes of the $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays can be written as
[…]
where $L_{\mu}=\bar{\nu_{\ell}}\gamma_{\mu}(1-\gamma_{5})\ell$ is the leptonic charged current, and $H^{\mu}=\langle D(k_{1})P(k_{2})|\bar{c}\gamma^{\mu}(1-\gamma_{5})\bar{b}|B(p_{B})\rangle$ is the hadronic matrix element. Usually, the hadronic matrix element $H^{\mu}$ can be obtained in terms of the form factors $F_{\bot,t,0,\|}$ of the $B\rightarrow DP$ transitions, which are similar to ones of $B\to PP$ transitions Boer:2016iez . Nevertheless, the calculations of the $F_{0},F_{t},F_{\perp},F_{\parallel}$ form factors are very difficult. In this work, the hadronic matrix element will be related by the SU(3) flavor symmetry.
Since the leptonic charged current $L_{\mu}$ is invariant under the SU(3) flavor symmetry, the hadronic matrix element $H^{\mu}$ can be parameterized by the SU(3) flavor symmetry. Then the decay amplitudes of the non-resonant $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays in Eq. (23) can be transformed as
$\mathcal{A}(B\rightarrow DP\ell^{+}\nu_{\ell})_{N}$ $=$ $\frac{G_{F}}{\sqrt{2}}V_{cb}^{*}H(B\rightarrow DP)_{N}\leavevmode\nobreak\ \bar{\nu}_{\ell}\gamma_{\mu}(1-\gamma_{5})\ell,$ (24)
@@ II.2 Decay amplitudes
where $c_{01,02}$ are the nonperturbative coefficients under the SU(3) flavor symmetry. The $c_{02}$ term is suppressed by the Okubo-Zweig-Iizuka (OZI) rule Okubo:1963fa ; Lipkin:1986bi ; Lipkin:1996ny , and it only appears in the decays with $\eta,\eta^{\prime}$ final states.
The idiographic hadronic amplitudes of the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays are given in Tab. 1. From Tab. 1, one can see that, if ignoring the OZI suppressed $c_{02}$ term, all hadronic amplitudes can be related by only one nonperturbative coefficient $c_{01}$.
The differential branching ratios of the nonresonant $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays are Boer:2016iez
with
where $\tau_{M}$($m_{M}$) is lifetime(mass) of $M$ particle. The ranges of integration are given by $(m_{D}+m_{P})^{2}\leq k^{2}\leq(m_{B}-m_{\ell})^{2}$ and $m_{\ell}^{2}\leq q^{2}\leq(m_{B}-\sqrt{k^{2}})^{2}$. If we ignore $|F_{t}|^{2}$ term since it is proportional to $m_{\ell}^{2}$ and it is small when $\ell=e,\mu$, $|H_{N}|^{2}$ is only include the hadronic part. Noted that although $|F_{t}|^{2}$ term might be large when $\ell=\tau$, it is difficult to estimate its contribution in this work, so we still ignore it. Then $H_{N}$, which only includes hadronic part, follow the relationship of the SU(3) flavor symmetry in the Tab. 1.
Table 1: SU(3) IRA amplitudes for $B\to DP\ell^{+}\nu$ decays due to $\bar{b}\to\bar{c}\ell^{+}\nu$.
   Decay modes Non-resonant decay amplitudes    Decay modes Non-resonant decay amplitudes
@@
II.3 Numerical results
@@
For the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, no any process has been measured until now. However, as given in Eq. (7), $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ has been measured. The $B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}}$ mode can decay via the non-resonance and the $D^{*}_{2}$ tensor meson resonance. In the later analysis of Sec. III, the contributions of $D^{*}_{2}$ tensor meson resonance are far less than the experimental data given in Eq. (7). So we think that the non-resonant contributions are dominant in the $B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}}$ decays, $i.e.$, $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}\approx\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$. The experimental data of $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ are used to determine $c_{01}$ in the non-resonant $B\to DP\ell^{+}\nu_{\ell^{\prime}}$ decays (Due to poor relevant experimental data, the OZI suppressed $c_{02}$ term is ignored). Then many other branching ratios of the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays can be predicted by using the data of $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$, which are listed in the second column of Table 2.
@@
From the second column of Table 2, one can see that many branching ratio central values of the non-resonant $B\to DP\ell^{\prime+}\nu_{\ell^{\prime}}$ decays, such as $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to\overline{D}^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}}$ $\mathcal{B}(B^{0}\to D^{-}_{s}K^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}_{s}\to\overline{D}^{0}K^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ $\mathcal{B}(B^{0}_{s}\to D^{-}\overline{K}^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ and $\mathcal{B}(B^{0}_{s}\to D_{s}^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, are on the orders of $10^{-4}$, which could be measured by the LHCb and Belle II experiments. Nevertheless, other decays, for examples, the non-resonant $B^{+}\to\overline{D}^{0}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$, $B^{0}\to D^{-}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$, $B^{0}_{s}\to D_{s}^{-}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$, and all $B\to DP\tau^{+}\nu_{\tau}$ decays, are strongly suppressed by the narrow phase spaces, their branching ratio central values are on the orders of $\mathcal{O}(10^{-5}-10^{-7})$, and they might not be observed by the experiments in the near future.
@@
Table 2: Branching ratios for $B\to DP\ell^{+}\nu_{\ell}$ decays due to $\bar{b}\to\bar{c}\ell^{+}\nu_{\ell}$ within $2\sigma$ errors. The unit is $10^{-3}$ for the branching ratios. $\mathcal{B}_{N}$ denotes the resonant branching ratios, $\mathcal{B}_{[R]}$ denotes the $R$ resonant branching ratios, edenotes experimental data within $2\sigma$ errors, and denotes the results obtained by considering the resonance width effects.

### v2

@@ Abstract
Decays $B\to DP\ell^{+}\nu_{\ell}\leavevmode\nobreak\ (\ell=e,\mu,\tau)$ with the non-resonance, the charmed vector resonances, the charmed scalar resonances and the charmed tensor resonances are explored by using the SU(3) flavor symmetry approach. Firstly, the decay amplitudes of different modes are related by the SU(3) flavor symmetry. Then, relevant experimental data are used to constrain non-perturbative coefficients in the non-resonant and various resonant $B\to DP\ell^{+}\nu_{\ell}$ decays. Finally, using the constrained non-perturbative coefficients, the branching ratios of not-yet-measured $B\to DP\ell^{+}\nu_{\ell}$ decays with the non-resonant and various charmed resonant contributions are predicted. Many branching ratios are predicted for the first time. We find that $B\to D\eta^{\prime}\ell^{+}\nu_{\ell},B_{s}\to D_{s}\eta^{\prime}\ell^{+}\nu_{\ell}$ decays only receive the non-resonant contributions, $B\to D_{s}K\ell^{+}\nu_{\ell}$, $B_{s}\to DK\ell^{+}\nu_{\ell}$, $B\to D\eta\ell^{+}\nu_{\ell}$ and $B_{s}\to D_{s}\eta\ell^{+}\nu_{\ell}$ decays receive both the non-resonant and the tensor resonant contributions, $B^{+}\to D^{-}\pi^{+}\ell^{+}\nu_{\ell}$ decays receive the non-resonant, the scalar resonant and the tensor resonant contributions, and other $B\to D\pi\ell^{+}\nu_{\ell}$ decays receive all kinds of contributions. These results can be used to test the SU(3) flavor symmetry approach in the four-body semileptonic $B$ decays in future experiments at LHCb and Belle-II.
@@ I Introduction
where $\mathcal{B}_{T,M}$ denote the total and $M$ resonant branching ratios. Note that, for $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ and $\mathcal{B}(B^{0}\to D^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ given in Eq. (1) and Eq. (4), the PDG reports results only include the $D_{0}$ and $D^{*}_{2}$ resonances, but do not include $D^{*}$ resonances. Present measurements of the $B\to DP\ell^{\prime+}\nu_{\ell^{\prime}}$ decays give us an opportunity to test theoretical approaches of the $B\to DP\ell^{+}\nu_{\ell}$ decays and to predict many non-measured decays, which can be further tested in the near future at LHCb and Belle-II.
@@
Theoretically, semileptonic decays are relatively simple, since the weak and strong dynamics can be separated in these decays. All the strong dynamics in the initial and final hadrons is included in the hadronic transition form factors. The calculations of the $B\to DP$ form factors are more complex than ones of the $B\to D$ form factors or the $B\to P$ form factors Feldmann:2018kqr . So the evaluations of the $B\to DP$ form factors are difficult. In the absence of reliable calculations, the symmetry analysis can provide very useful information about the decays. SU(3) flavor symmetry is one of the popular symmetry approaches. And it has been widely used to study $b$-hadron decays He:1998rq ; He:2000ys ; Fu:2003fy ; Hsiao:2015iiu ; He:2015fwa ; He:2015fsa ; Deshpande:1994ii ; Gronau:1994rj ; Gronau:1995hm ; Shivashankara:2015cta ; Zhou:2016jkv ; Cheng:2014rfa ; Wang:2021uzi ; Wang:2020wxn , $c$-hadron decays Wang:2021uzi ; Wang:2020wxn ; Grossman:2012ry ; Pirtskhalava:2011va ; Savage:1989qr ; Savage:1991wu ; Altarelli:1975ye ; Lu:2016ogy ; Geng:2017esc ; Geng:2018plk ; Geng:2017mxn ; Geng:2019bfz ; Wang:2017azm ; Wang:2019dls ; Wang:2017gxe ; Qiao:2024srw ; Muller:2015lua , and light hadron decays Wang:2019alu ; Wang:2021uzi ; Xu:2020jfr ; Chang:2014iba ; Zenczykowski:2005cs ; Zenczykowski:2006se ; Cabibbo:1963yz . SU(3) flavor symmetry breaking effects due to the mass differences between the $u,d$ and $s$ quarks have also been studied, for instance, in Refs. Martinelli:2022xir ; Wang:2022fbk ; Imbeault:2011jz ; Wu:2005hi ; Dery:2020lbc ; Sasaki:2008ha ; Pham:2012db ; Geng:2018bow ; Flores-Mendieta:1998tfv ; Cheng:2012xb ; Xu:2013dta ; He:2014xha ; Yang:2015era .
@@
Some four-body semileptonic decays $B/D\to PP\ell^{+}\nu_{\ell}$ and $B\to D^{(*)}P\ell^{+}\nu_{\ell}$ have been investigated Cheng:1993ah ; Tsai:2021ota ; Feldmann:2018kqr ; Kim:1999gm ; Kim:1998nn ; LeYaouanc:2018zyo ; Boer:2016iez ; Hambrock:2015aor ; Faller:2013dwa ; Shi:2021bvy ; Achasov:2020qfx ; Wiss:2007mr ; Wang:2016wpc ; Achasov:2021dvt ; Gustafson:2023lrz . In this work, we will explore the $B\to DP\ell^{+}\nu_{\ell}$ decays with the non-resonant, the charmed vector resonant, the charmed scalar resonant and the charmed tensor resonant contributions by the SU(3) flavor symmetry based on the experimental data. Firstly, the hadronic amplitude relations or the form factor relations between different decay modes will be constructed. Then, the hadronic amplitudes or the form factors will be extracted by using the available data. Finally, the not-yet-measured modes will be predicted for further tests in experiments.
@@
This paper is organized as follows. The non-resonant contributions of the $B\to DP\ell\nu_{\ell}$ decays are discussed in Sec. II. The charmed vector resonant, the charmed scalar resonant and the charmed tensor resonant contributions of the $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays are presented in Sec. III. Finally, summary is given in Sec. IV.
@@ II Non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays
II.1 Theoretical framework
The non-resonant $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays are generated by $\bar{b}\rightarrow\bar{c}\ell^{+}\nu_{\ell}$ transition, and the effective Hamiltonian is
$\mathcal{H}_{eff}(\bar{b}\rightarrow\bar{c}\ell^{+}\nu_{\ell})=\frac{G_{F}}{\sqrt{2}}V_{cb}\bar{c}\gamma^{\mu}(1-\gamma_{5})b\leavevmode\nobreak\ \bar{\nu_{\ell}}\gamma_{\mu}(1-\gamma_{5})\ell,$ (8)
where $G_{F}$ is the Fermi constant, and $V_{cb}$ is the CKM matrix element. Decay amplitudes of the non-resonant $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays can be written as
[…]
where $L_{\mu}=\bar{\nu_{\ell}}\gamma_{\mu}(1-\gamma_{5})\ell$ is the leptonic charged current, and $H^{\mu}=\langle D(k_{1})P(k_{2})|\bar{c}\gamma^{\mu}(1-\gamma_{5})b|B(p_{B})\rangle$ is the hadronic amplitude. Usually, $H^{\mu}$ can be obtained in terms of the form factors $F_{\bot,t,0,\|}$ of the $B\rightarrow DP$ transitions, which are are non-perturbative objects and are similar to ones of $B\to PP$ transitions Boer:2016iez . Nevertheless, the calculations of the $F_{0},F_{t},F_{\perp},F_{\parallel}$ are very difficult. In this work, the hadronic amplitude will be related by the SU(3) flavor symmetry.
Relevant meson multiplets are given first. Bottom pseudoscalar triplet $B_{i}$, charm pseudoscalar triplet $D_{i}$, charm scalar triplet $D_{0i}$, charm vector triplet $D^{*}_{i}$ and charm tensor triplet $D^{*}_{2i}$ under the SU(3) flavor symmetry of the $u$, $d$, $s$ quarks are
@@ II.1 Theoretical framework
$D^{*}_{2i}$ $=$ $\Big{(}\overline{D}^{*0}_{2}(\bar{c}u),D^{*-}_{2}(\bar{c}d),D^{*-}_{s2}(\bar{c}s)\Big{)},$ (12)
where $i=1,2,3$ for $u$, $d$, $s$ quarks. Note that the structures of scalar $D_{0}$ mesons are not known well, and they might be four-quark states, $D_{0i}^{4q}=\Big{(}\overline{D}^{0}_{0}(\bar{c}u\bar{d}d),D^{-}_{0}(\bar{c}d\bar{u}u),D^{-}_{s0}(\bar{c}s(\bar{u}u+\bar{d}d)/\sqrt{2})\Big{)}$ Cheng:2003kg . Note that $D_{0i}$, $D^{*}_{i}$ and $D^{*}_{2i}$ will be used for the resonances in Sec. III, and they are given here together. Light pseudoscalar octets and singlets $P^{i}_{j}$ are
@@
The leptonic charged current is invariant under the SU(3) flavor symmetry, and the hadronic amplitude of the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decay can be parameterized by the SU(3) flavor symmetry as
@@
where $c_{01,02}$ are the non-perturbative coefficients under the SU(3) flavor symmetry. The $c_{02}$ term is suppressed by the Okubo-Zweig-Iizuka (OZI) rule Okubo:1963fa ; Lipkin:1986bi ; Lipkin:1996ny , and it only appears in the decays with $\eta,\eta^{\prime}$ final states. The idiographic hadronic amplitudes of the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays are given in Tab. 1. From Tab. 1, one can see that, if ignoring the OZI suppressed $c_{02}$ contributions, all hadronic amplitudes may be related by the nonperturbative coefficient $c_{01}$.
Table 1: The hadronic amplitudes for the non-resonant $B\to DP\ell^{+}\nu$ decays under the SU(3) flavor symmetry.
   Decay modes Hadronic amplitudes    Decay modes Hadronic amplitudes
@@
The differential branching ratios of the non-resonant $B\rightarrow DP\ell^{+}\nu_{\ell}$ decays are Boer:2016iez
with
where $\tau_{M}$($m_{M}$) is lifetime(mass) of $M$ particle. The ranges of integration are given by $(m_{D}+m_{P})^{2}\leq k^{2}\leq(m_{B}-m_{\ell})^{2}$ and $m_{\ell}^{2}\leq q^{2}\leq(m_{B}-\sqrt{k^{2}})^{2}$. If we ignore $|F_{t}|^{2}$ term since it is proportional to $m_{\ell}^{2}$ and it is small when $\ell=e,\mu$, $|H_{N}|^{2}$ is only include the hadronic part. Noted that although $|F_{t}|^{2}$ term might be large when $\ell=\tau$, it is difficult to estimate its contribution in this work, so we still ignore it. Then $H_{N}$, which only includes hadronic part, follow the relationship of the SU(3) flavor symmetry in Tab. 1.
II.2 Numerical results
@@
For the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, no any process has been measured until now. However, as given in Eq. (7), $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ has been measured. The $B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}}$ mode can decay via the non-resonance and the $D^{*}_{2}$ tensor meson resonance. In the subsequent analysis in Sec. III, the contributions of $D^{*}_{2}$ tensor meson resonance are far less than the experimental data given in Eq. (7). So we think that the non-resonant contributions are dominant in the $B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}}$ decays, $i.e.$, $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}\approx\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$. The experimental data of $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ are used to determine $c_{01}$ in the non-resonant $B\to DP\ell^{+}\nu_{\ell^{\prime}}$ decays (Due to poor relevant experimental data, the OZI suppressed $c_{02}$ term is ignored). Then many other branching ratios of the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays can be predicted by using the data of $\mathcal{B}(B^{+}\to D^{-}_{s}K^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$, which are listed in the second column of Tab. 2.
@@
From the second column of Tab. 2, one can see that many branching ratio central values of the non-resonant $B\to DP\ell^{\prime+}\nu_{\ell^{\prime}}$ decays, such as $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to\overline{D}^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}}$ $\mathcal{B}(B^{0}\to D^{-}_{s}K^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}_{s}\to\overline{D}^{0}K^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ $\mathcal{B}(B^{0}_{s}\to D^{-}\overline{K}^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ and $\mathcal{B}(B^{0}_{s}\to D_{s}^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, are on the orders of $10^{-4}$, which could be measured by the LHCb and Belle II experiments. Nevertheless, other decays, for example, the non-resonant $B^{+}\to\overline{D}^{0}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$, $B^{0}\to D^{-}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$, $B^{0}_{s}\to D_{s}^{-}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$, and all $B\to DP\tau^{+}\nu_{\tau}$ decays, are strongly suppressed by the narrow phase spaces, their branching ratio central values are on the orders of $\mathcal{O}(10^{-5}-10^{-7})$, and they might not be observed by the experiments in the near future.
@@
Table 2: Branching ratios for the $B\to DP\ell^{+}\nu_{\ell}$ decays within $2\sigma$ errors (in units of $10^{-3}$). $\mathcal{B}_{N}$ denotes the non-resonant branching ratios, $\mathcal{B}_{[R]}$ denotes the $R$ resonant ones, edenotes experimental data within $2\sigma$ errors, and denotes the results obtained by considering the resonance width effects.